Hydrogen produced from a hydrolysis reaction was collected over water. The total volume of h2 (g) was 93.26 at the temperature of 25 degree C. The barometric pressure was 742 mmHg and the vapor pressure of water at 25 degrees C was 24.0 mmHg. Calculate the moles of hydrogen gas produced by the reaction


1
Expert's answer
2021-02-03T07:15:42-0500

Convert mmHg to Pa: 742 mmHg is 98925.2 Pa, 24 mmHg is 3199.74 Pa. Assume the unit of volume is L (so that it is 93.26 L).


The barometric pressure is total pressure of the gas and water vapor:


"P_B=P_V+P_{\text{H}_2}\space\rightarrow\\
P_{\text{H}_2}=P_B-P_V."

Now apply the ideal gas law:


"PV=nRT,\\\space\\
n=\frac{PV}{RT}=\frac{(P_B-P_V)V}{RT},\\\space\\
n=\frac{(98925.2-3199.74)93.26\cdot10^{-3}}{8.314(273+25)}=3.6\text{ mol}."